How much is the length of Cs+ cation when Chloride ion is 1.81 Armstrong?(Cell parameter a=4.124 Armstrong)
Answer #1

Answer)

CsCl structure type is bcc.

Cl- anion is located at the corner of the lattice

Cs+ cation located at the center of the lattice

Cell parameter (a) =2/ 3 (radius of Cs+ + radius of Cl-) = 4.124

radius of Cl- = 1.81

(radius of Cs+ + 1.81 ) = 4.124 x 3 /2 = 3.571

radius of Cs+ = 3.571 - 1.81 = 1.761

..........................................

​​​​​​​If you have found it useful please give a like .........